# Cosmic Ray Muons in Laboratories Deep Underground

### v1

@@ I Introduction
For many decades, measurements of muons have been conducted underwater and underground to study the energy spectrum and composition of primary cosmic rays Barrett et al. (1952); Workman et al. (2022). Despite these long-standing efforts, large uncertainties surrounding atmospheric muon and high-energy neutrino fluxes persist, with sustained challenges to their reduction Barr et al. (2006); Honda et al. (2007); Yañez and Fedynitch (2023). Shielded by the Earth’s natural overburden, large-volume underground detectors offer a unique opportunity to study high-energy muons without the technical challenges or high cost of single-purpose surface-based spectrometers capable of resolving multi-hundred TeV range energies.
@@
Addressing these uncertainties is of paramount importance for several reasons. Precise estimation of the muon flux can significantly benefit fields like rare-event searches, including direct dark matter searches. The detectors for these searches are typically installed deep underground in mines or beneath mountains in order to limit exposure to cosmic-ray muons Formaggio and Martoff (2004). However, cosmic-ray muons produce secondary neutrons by interacting with the rock or surrounding detector materials, which pose significant identification challenges, as they can mimic low-energy dark matter signals. Knowledge of the underground muon rate, energy spectrum, and angular distribution is necessary to estimate muon-induced backgrounds to design effective shielding, including muon and neutron vetos. For the past few decades, analytical calculations Bugaev et al. (1998); Lipari and Stanev (1991) and empirical parametric fits to vertical-equivalent muon intensity data Crouch (1987); Mei and Hime (2006) have served as inputs to Monte Carlo muon transport codes such as mum Sokalski et al. (2001), music and musun Kudryavtsev (2009), and Geant4 Agostinelli et al. (2003). However, these methods often demand significant computing resources and do not thoroughly address the treatment of errors. fluka Battistoni et al. (2015), meanwhile, can be used to calculate muon fluxes starting from a cosmic ray flux parameterization; however, it also comes with computational cost, in particular for high energies.
@@
In our previous work in Ref. Fedynitch et al. (2022), we concentrated on the development of the open-source mute code, designed to calculate underground muon fluxes and other related observables without using those fluxes as input. mute is flexible, efficient, and precise, and combines two modern computational tools, mceq Fedynitch (2015); Fedynitch et al. (2019) and proposal Koehne et al. (2013), for surface muon fluxes and underground transport respectively. We previously compared vertical-equivalent underground intensities computed with mute using a flat overburden to experimental measurements. In the present study, we expand on this by scrutinizing the residual body of related data using an updated version of mute in combination with the latest and most precise surface flux model, daemonflux model Yañez and Fedynitch (2023). daemonflux makes use of two data-driven models to achieve uncertainties on surface fluxes of less than 10% up to 1 TeV: Global Spline Fit (gsfDembinski et al. (2018) for the primary cosmic ray flux, and the Data-Driven Model (ddmFedynitch and Huber (2022) for hadronic interactions. In this work, we calculate total underground and underwater muon fluxes, as well as underground angular distributions, energy spectra, and mean energies. We use topographic maps for laboratories situated under mountains and we incorporate the density and composition of the rock above each lab into our calculations, which are significant sources of systematic uncertainty. We compare our mute predictions with experimental data from different underground and underwater sites worldwide. The new release of mute is publicly available so our results can be easily reproduced and used for further studies.111mute v2 is publicly available at https://github.com/wjwoodley/mute and v3 will be made publicly available soon.
@@ II Computational Method
For practical applications, the muon intensity at a certain depth underground has been historically parameterized using functions termed Depth-Intensity Relations (DIR). D.-M. Mei and A. Hime Mei and Hime (2006) give the following DIR for depth ranges between 1 and 10 km.w.e.:
@@
where $A$, $h_{0}$, and $\alpha$ are fit from data. Parametric fits like these can provide reasonable estimates for underground labs, as demonstrated in Ref. Mei and Hime (2006). There are, however, several caveats. Firstly, the fits are done to vertical-equivalent underground intensity data, which is known to only approximately correspond to the true vertical intensity at zenith angles below $\sim 30^{\circ}$ Barrett et al. (1952); Bugaev et al. (1998); Fedynitch et al. (2022). In addition, various data sets are typically combined without compensating for systematic uncertainties through correction functions or nuisance parameters, and hence the fit results might be significantly biased towards experiments that have underestimated their uncertainties. They are also limited in the sense that they cannot provide additional physical information about the muons underground, such as their energy and angular distributions, which must be calculated using separate parameterizations.
@@
In this follow-up study, we introduce mute v2 Woodley and Fedynitch (2022) and v3, capable of calculating muon observables for laboratories beneath mountains. For the case of a flat geometry above the lab (labs under flat overburdens or underwater), due to the assumed symmetry in the azimuthal angle, Eq. 5 provides a direct geometric correspondence between slant depth and zenith angle for a given vertical depth. This relationship reduces the underground or underwater flux in Eq. 3 at the specified depth to a function of just two variables: $\Phi^{u}(E^{u},\theta)$. For mountains, however, the non-uniform shape of the overburden introduces dependence on the azimuthal angle to the amount of rock a muon has to travel through, meaning the geometry of the mountain has to be taken into account in the calculations. To do this, topographic maps of the mountains in terms of the depth in spherical coordinates, $X(\theta,\phi)$, are obtained from the labs, such as the one for Laboratori Nazionali del Gran Sasso (LNGS) in Italy, shown in Fig. 1. For this reason, slant depth and zenith angle are kept as separate variables in the calculation of underground fluxes for mountains: $\Phi^{u}(E^{u},X(\theta,\phi),\theta)$. The explicit dependence on $\theta$ is relevant because the surface fluxes depend on the zenith angle but not the azimuthal angle due to the symmetry at high energies.
Figure 1: Slant depths of the Gran Sasso mountain, $X(\theta,\phi)$, in terms of zenith and azimuthal angle Aglietta et al. (1998). Depths greater than 14 km.w.e. have been masked (shown in black), as 14 km.w.e. is the default maximum slant depth for calculations by mute. In all cases, muons traveling along $\theta=0^{\circ}$ are defined as vertical down-going muons.
@@
Using these underground intensities, we calculate various physical observables for different underground sites worldwide. The underground sites used are listed in Table 1, along with their average rock density and vertical depths in km as found in the literature. We present flat overburden calculations for the Waste Isolation Pilot Plant (WIPP), Soudan Underground Laboratory, Boulby Underground Laboratory, Stawell Underground Physics Laboratory (SUPL), Sanford Underground Research Facility (SURF) located in the former Homestake Gold Mine, and the Sudbury Neutrino Observatory Laboratory (SNOLAB). For laboratories under mountains, we have used topographic maps of the Yangyang Underground Laboratory (Y2L) centered around the COSINE-100 detector, the Kamioka Observatory under the Ikenoyama mountain centered around two locations — the KamLAND Abe et al. (2010) and Super-Kamiokande Abe et al. (2010); Fukuda et al. (2003) detectors — LNGS under the Gran Sasso mountain centered around the Large Volume Detector (LVD) Aglietta et al. (1998), Laboratoire Souterrain de Modane (LSM) centered around the Fréjus detector Berger et al. (1989), and China Jinping Underground Laboratory (CJPL-I) centered around the Jinping Neutrino Experiment (JNE) Guo et al. (2021). In the case of labs under mountains, it is necessary to center the map around a specific detector’s location due to the calculations’ dependence on azimuthal angle. We selected these labs based on the availability of experimental data with which we can compare our results.
@@ III.2 Rock Composition
The effect of changing the $Z/A\right>$ and $Z^{2}/A\right>$ values on the energy loss results from proposal can be significant. Figure 3 shows the ratio of underground muon intensities using the laboratory rock from Table 2 to those using standard rock, demonstrating that increasing the $Z^{2}/A\right>$ value leads to a decrease in the muon intensity results, as physically expected. There is a deviation from standard rock of -27% in the underground intensity for CJPL-I at the vertical depth of the lab (approximately 6 km.w.e.), while the effect is minimal for Y2L, Kamioka, and LNGS at their respective vertical depths (less than -5%). However, the effect is amplified as slant depth increases, as the muons travel through higher quantities of rock. Therefore, muons reaching the labs at higher zenith angles are affected more greatly by changes to the chemical composition. Because of this depth-dependent energy loss effect, DIRs and mean underground energy parameterizations that are not corrected for rock type, like those in Eqs. 1 and 2, are less suitable for high-precision calculations, highlighting the need for detailed simulations in underground muon flux calculations.
@@ III.3 Sternheimer Parameters
To describe ionization losses, proposal uses the Bethe-Bloch equation Workman et al. (2022). This equation includes a density correction term, $-\delta/2$, as described in Refs. Sternheimer (1952); Koehne et al. (2013), where $\delta$ is given by
@@ IV.1 Total Flux Underground
The total underground flux has been calculated for the sites listed in Table 1 using daemonflux as the surface muon flux model and U.S. Standard Atmosphere atm (1976) as the atmospheric density model. The results are given in Table 3 and are shown in Fig. 4 along with experimental measurements.
@@
The main uncertainties in mute come from the surface flux model, daemonflux, and the modeling of the overburden. The former are represented by the gray error bands in Figs. 4 and 5, and contribute an error of $\sim$7for deep depths (higher muon energies) and around 2% for shallow depths. The latter are represented by horizontal gray bars and are entirely attributed to the range of allowed average rock densities listed in Table 1. These two errors are shown separately to emphasize their independence from each other. The interaction model uncertainties are independent of the lab, in contrast to the knowledge of the rock density, which is either limited by the precision of the geological surveys or by the non-uniformity of the overburden, making it difficult to compute a representative mean density. The interaction model errors have been discussed explicitly in Ref. Fedynitch et al. (2022), where the rock density variations were implicitly included in the errors of the experimental data. The magnitudes of the uncertainties in Table 1 are large in many cases, leading to significant variations in the total flux seen in Fig. 5. Error bars are not shown for labs where no uncertainty on the rock density was found in the literature. The mountain maps cited in Section II are assumed to be exact, so no error was considered for the slant depth values. Lastly, the proposal Monte Carlo simulations in mute were performed with $N=10^{6}$ events per energy and zenith angle bin, making the statistical error across all calculations negligible.
@@ IV.1.3 SUPL
As stated in Section III, aside from the rock density, no additional relevant quantitative information exists about the rock above SUPL. For this reason, standard rock was used in the proposal simulations. Despite not considering the rock composition, the mute result agrees well with the SABRE measurement from Ref. Melbourne (2021). We note that the result may change with more information about the rock density and composition.
@@ IV.1.5 SNOLAB
Although agreement can be achieved between the mute prediction and the SNO data point from Ref. Aharmim et al. (2009) when all systematic uncertainties are considered, it should be noted that this data point is not well-understood, and, like for other laboratories, agreement is dependent on the depth we consider for the lab. Despite some ambiguity about the correct depth to use for SNOLAB, in our calculations, we take $h=5.920$ km.w.e. based on the depth $d=2.092$ km quoted in Ref. Aharmim et al. (2009).
@@ IV.2 Total Flux Underwater
When performing computations for laboratories underwater as opposed to under a flat overburden, the only change in the mute calculation is to the medium used for the surface-to-underground transfer tensor generation with proposal, from rock to water. Thus, the accuracy of the mute predictions in Fig. 5 seen for rock should apply to water as well. We note that KM3NeT’s data is a factor of 2 below our prediction for sea water. A more recent measurement of the event rate as a function of zenith angle from a more complete version of the KM3NeT detector Aiello et al. (2024) indicates a 30–40% discrepancy to a calculation made using a similar technique and the sibyll-2.3d interaction model. Their result is approximately in agreement with our findings illustrated in Fig. 11 (see Appendix C) and also confirms that the previous results from Ref. Ageron et al. (2020) shown in Fig. 6 were significantly underestimated.
@@ V Angular Distributions
Figure 7: Underground double-differential muon intensities for the Gran Sasso mountain as calculated by mute (top) and the residuals with respect to the LVD data (bottom) as a function of zenith and azimuthal angle. In the figure of the residuals, black indicates masked bins for which the slant depth is greater than 14 km.w.e. Additionally, because the binning on the map of the Gran Sasso mountain is very fine, with large bin-to-bin fluctuations in the residuals, a Gaussian filter ($\sigma=1.5$) has been applied to help identify any characteristic trends.
The intensities were similarly computed using mute with the method described in Section II, and the results are shown in Fig. 7 (top). In order to compare the experimental and predicted intensities, the residuals for each $(\theta,\phi)$-bin of the data were calculated as
@@
where $\delta I^{u}_{\mathrm{LVD}}$ are the statistical errors in each angular bin. The result is shown in Fig. 7 (bottom). Since, by construction from Eq. 11, the normalizations of the mute result and the LVD data match, only relative trends can be observed. The calculation is slightly below the data at near-vertical directions and above it for intermediate zenith angles, except for a small patch in the SW direction ($\sim 225^{\circ}$ azimuth), where the situation is inverted. We calculate the $\chi^{2}/\mathrm{ndf}$ to be 1.34, where $\mathrm{ndf}$ is assumed to be the number of non-zero bins ($\mathrm{ndf}=31756$).
@@
In general, comparing the experimental and calculated angular distributions can reveal critical information about the accuracy of data analyses of underground experiments. The overall agreement obtained with the LVD experimental data suggests that mute can be a powerful and helpful tool to inform on potential sources of errors and mis-reconstructions in data analyses for underground experiments, mainly due to the absence of statistical errors.

### v2

@@ I Introduction
For many decades, measurements of muons have been conducted underwater and underground to study the energy spectrum and composition of primary cosmic rays Barrett et al. (1952); Workman et al. (2022). Despite these long-standing efforts, large uncertainties surrounding atmospheric muon and high-energy neutrino fluxes persist, with sustained challenges to their reduction Barr et al. (2006); Honda et al. (2007); Yañez and Fedynitch (2023). Shielded by the Earth’s natural overburden, large-volume underground detectors offer a unique opportunity to study high-energy muons without the technical challenges or high cost of single-purpose surface-based spectrometers capable of resolving multihundred TeV range energies.
@@
Addressing these uncertainties is of paramount importance for several reasons. Precise estimation of the muon flux can significantly benefit fields like rare-event searches, including direct dark matter searches. The detectors for these searches are typically installed deep underground in mines or beneath mountains in order to limit exposure to cosmic-ray muons Formaggio and Martoff (2004). However, cosmic-ray muons produce secondary neutrons by interacting with the rock or surrounding detector materials, which pose significant identification challenges, as they can mimic low-energy dark matter signals. Knowledge of the underground muon rate, energy spectrum, and angular distribution is necessary to estimate muon-induced backgrounds to design effective shielding, including muon and neutron vetos. For the past few decades, analytical calculations Bugaev et al. (1998); Lipari and Stanev (1991) and empirical parametric fits to vertical-equivalent muon intensity data Crouch (1987); Mei and Hime (2006) have served as inputs to Monte Carlo muon transport codes such as mum Sokalski et al. (2001), music and musun Kudryavtsev (2009), and Geant4 Agostinelli et al. (2003). However, these methods often demand significant computing resources and do not thoroughly address the treatment of errors. fluka Battistoni et al. (2015), meanwhile, can be used to calculate muon fluxes starting from a cosmic ray flux parametrization; however, it also comes with computational cost, in particular for high energies.
@@
In our previous work in Ref. Fedynitch et al. (2022), we concentrated on the development of the open-source mute code, designed to calculate underground muon fluxes and other related observables without using those fluxes as input. mute is flexible, efficient, and precise, and combines two modern computational tools, mceq Fedynitch (2015); Fedynitch et al. (2019) and proposal Koehne et al. (2013), for surface muon fluxes and underground transport respectively. We previously compared vertical-equivalent underground intensities computed with mute using a flat overburden to experimental measurements. In the present study, we expand on this by scrutinizing the residual body of related data using an updated version of mute in combination with the latest and most precise surface flux model, the daemonflux model Yañez and Fedynitch (2023). daemonflux makes use of two data-driven models to achieve uncertainties on surface fluxes of less than 10% up to 1 TeV: Global Spline Fit (gsfDembinski et al. (2018) for the primary cosmic ray flux, and the Data-Driven Model (ddmFedynitch and Huber (2022) for hadronic interactions. In this work, we calculate total underground and underwater muon fluxes, as well as underground angular distributions, energy spectra, and mean energies. We use topographic maps for laboratories situated under mountains and we incorporate the density and composition of the rock above each lab into our calculations, which are significant sources of systematic uncertainty. We compare our mute predictions with experimental data from different underground and underwater sites worldwide. The new release of mute is publicly available so our results can be easily reproduced and used for further studies. [^1] https://github.com/wjwoodley/mute
@@ II Computational Method
For practical applications, the muon intensity at a certain depth underground has been historically parametrized using functions termed depth-intensity relations (DIR). D.-M. Mei and A. Hime Mei and Hime (2006) give the following DIR for depth ranges between 1 and 10 km.w.e.:
@@
where $A$, $h_{0}$, and $\alpha$ are fit from data. Parametric fits like these can provide reasonable estimates for underground labs, as demonstrated in Ref. Mei and Hime (2006). There are, however, several caveats. Firstly, the fits are done to vertical-equivalent underground intensity data, which is known to only approximately correspond to the true vertical intensity at zenith angles below $\sim 30^{\circ}$ Barrett et al. (1952); Bugaev et al. (1998); Fedynitch et al. (2022). In addition, various datasets are typically combined without compensating for systematic uncertainties through correction functions or nuisance parameters, and hence the fit results might be significantly biased towards experiments that have underestimated their uncertainties. They are also limited in the sense that they cannot provide additional physical information about the muons underground, such as their energy and angular distributions, which must be calculated using separate parametrizations.
@@
In this followup study, we introduce mute v2 Woodley and Fedynitch (2022) and v3, capable of calculating muon observables for laboratories beneath mountains. For the case of a flat geometry above the lab (labs under flat overburdens or underwater), due to the assumed symmetry in the azimuthal angle, Eq. 5 provides a direct geometric correspondence between the slant depth and zenith angle for a given vertical depth. This relationship reduces the underground or underwater flux in Eq. 3 at the specified depth to a function of just two variables: $\Phi^{u}(E^{u},\theta)$. For mountains, however, the nonuniform shape of the overburden introduces dependence on the azimuthal angle to the amount of rock a muon has to travel through, meaning the geometry of the mountain has to be taken into account in the calculations. To do this, topographic maps of the mountains in terms of the depth in spherical coordinates, $X(\theta,\phi)$, are obtained from the labs, such as the one for Laboratori Nazionali del Gran Sasso (LNGS) in Italy, shown in Fig. 1. For this reason, the slant depth and the zenith angle are kept as separate variables in the calculation of underground fluxes for mountains: $\Phi^{u}(E^{u},X(\theta,\phi),\theta)$. The explicit dependence on $\theta$ is relevant because the surface fluxes depend on the zenith angle but not the azimuthal angle due to the symmetry at high energies.
Figure 1: Slant depths of the Gran Sasso mountain, $X(\theta,\phi)$, in terms of the zenith and azimuthal angles Aglietta et al. (1998). Depths greater than 14 km.w.e. have been masked (shown in black), as 14 km.w.e. is the default maximum slant depth for calculations by mute. In all cases, muons traveling along $\theta=0^{\circ}$ are defined as vertical down-going muons.
@@
Using these underground intensities, we calculate various physical observables for different underground sites worldwide. The underground sites used are listed in Table 1, along with their average rock density and vertical depths in km as found in the literature. We present flat overburden calculations for the Waste Isolation Pilot Plant (WIPP), Soudan Underground Laboratory, Boulby Underground Laboratory, Stawell Underground Physics Laboratory (SUPL), Sanford Underground Research Facility (SURF) located in the former Homestake Gold Mine, and the Sudbury Neutrino Observatory Laboratory (SNOLAB). For laboratories under mountains, we have used topographic maps of the Yangyang Underground Laboratory (Y2L) centered around the COSINE-100 detector, the Kamioka Observatory under the Ikenoyama mountain centered around two locations—the KamLAND Abe et al. (2010) and Super-Kamiokande Abe et al. (2010); Fukuda et al. (2003) detectors—LNGS under the Gran Sasso mountain centered around the Large Volume Detector (LVD) Aglietta et al. (1998), Laboratoire Souterrain de Modane (LSM) centered around the Fréjus detector Berger et al. (1989), and China Jinping Underground Laboratory (CJPL-I) centered around the Jinping Neutrino Experiment (JNE) Guo et al. (2021). In the case of labs under mountains, it is necessary to center the map around a specific detector’s location due to the calculations’ dependence on the azimuthal angle. We selected these labs based on the availability of experimental data with which we can compare our results.
@@ III.2 Rock composition
The effect of changing the $Z/A\right>$ and $Z^{2}/A\right>$ values on the energy loss results from proposal can be significant. Figure 3 shows the ratio of underground muon intensities using the laboratory rock from Table 2 to those using standard rock, demonstrating that increasing the $Z^{2}/A\right>$ value leads to a decrease in the muon intensity results, as physically expected. There is a deviation from standard rock of -27% in the underground intensity for CJPL-I at the vertical depth of the lab (approximately 6 km.w.e.), while the effect is minimal for Y2L, Kamioka, and LNGS at their respective vertical depths (less than -5%). However, the effect is amplified as slant depth increases, as the muons travel through higher quantities of rock. Therefore, muons reaching the labs at higher zenith angles are affected more greatly by changes to the chemical composition. Because of this depth-dependent energy loss effect, DIRs and mean underground energy parametrizations that are not corrected for rock type, like those in Eqs. 1 and 2, are less suitable for high-precision calculations, highlighting the need for detailed simulations in underground muon flux calculations.
@@ III.3 Sternheimer parameters
To describe ionization losses, proposal uses the Bethe–Bloch equation Workman et al. (2022). This equation includes a density correction term, $-\delta/2$, as described in Refs. Sternheimer (1952); Koehne et al. (2013), where $\delta$ is given by
@@ IV.1 Total flux underground
The total underground flux has been calculated for the sites listed in Table 1 using daemonflux as the surface muon flux model and U.S. Standard Atmosphere on Extension to the Standard Atmosphere (1976) as the atmospheric density model. The results are given in Table 3 and are shown in Fig. 4 along with experimental measurements.
@@
The main uncertainties in mute come from the surface flux model, daemonflux, and the modeling of the overburden. The former are represented by the gray error bands in Figs. 4 and 5, and contribute an error of $\sim$7for deep depths (higher muon energies) and around 2% for shallow depths. The latter are represented by horizontal gray bars and are entirely attributed to the range of allowed average rock densities listed in Table 1. These two errors are shown separately to emphasize their independence from each other. The interaction model uncertainties are independent of the lab, in contrast to the knowledge of the rock density, which is either limited by the precision of the geological surveys or by the nonuniformity of the overburden, making it difficult to compute a representative mean density. The interaction model errors have been discussed explicitly in Ref. Fedynitch et al. (2022), where the rock density variations were implicitly included in the errors of the experimental data. The magnitudes of the uncertainties in Table 1 are large in many cases, leading to significant variations in the total flux seen in Fig. 5. Error bars are not shown for labs where no uncertainty on the rock density was found in the literature. The mountain maps cited in Sec. II are assumed to be exact, so no error was considered for the slant depth values. Lastly, the proposal Monte Carlo simulations in mute were performed with $N=10^{6}$ events per energy and zenith angle bin, making the statistical error across all calculations negligible.
@@ IV.1.3 SUPL
As stated in Sec. III, aside from the rock density, no additional relevant quantitative information exists about the rock above SUPL. For this reason, standard rock was used in the proposal simulations. Despite not considering the rock composition, the mute result agrees well with the SABRE measurement from Ref. Melbourne (2021). We note that the result may change with more information about the rock density and composition.
@@ IV.1.5 SNOLAB
Although agreement can be achieved between the mute prediction and the SNO data point from Ref. Aharmim et al. (2009) when all systematic uncertainties are considered, it should be noted that this data point is not well understood, and, like for other laboratories, agreement is dependent on the depth we consider for the lab. Despite some ambiguity about the correct depth to use for SNOLAB, in our calculations, we take $h=5.920$ km.w.e. based on the depth $d=2.092$ km quoted in Ref. Aharmim et al. (2009).
@@ IV.2 Total flux underwater
When performing computations for laboratories underwater as opposed to under a flat overburden, the only change in the mute calculation is to the medium used for the surface-to-underground transfer tensor generation with proposal, from rock to water. Thus, the accuracy of the mute predictions in Fig. 5 seen for rock should apply to water as well. We note that KM3NeT’s data is a factor of 2 below our prediction for sea water. A more recent measurement of the event rate as a function of the zenith angle from a more complete version of the KM3NeT detector Aiello et al. (2024) indicates a 30–40% discrepancy to a calculation made using a similar technique and the sibyll-2.3d interaction model. Their result is approximately in agreement with our findings illustrated in Fig. 11 (see Appendix C) and also confirms that the previous results from Ref. Ageron et al. (2020) shown in Fig. 6 were significantly underestimated.
@@ V Angular Distributions
Figure 7: Underground double-differential muon intensities for the Gran Sasso mountain as calculated by mute (top) and the residuals with respect to the LVD data (bottom) as a function of the zenith and azimuthal angles. In the figure of the residuals, black indicates masked bins for which the slant depth is greater than 14 km.w.e. Additionally, because the binning on the map of the Gran Sasso mountain is very fine, with large bin-to-bin fluctuations in the residuals, a Gaussian filter ($\sigma=1.5$) has been applied to help identify any characteristic trends.
The intensities were similarly computed using mute with the method described in Sec. II, and the results are shown in Fig. 7 (top). In order to compare the experimental and predicted intensities, the residuals for each $(\theta,\phi)$-bin of the data were calculated as
@@
where $\delta I^{u}_{\mathrm{LVD}}$ are the statistical errors in each angular bin. The result is shown in Fig. 7 (bottom). Since, by construction from Eq. 11, the normalizations of the mute result and the LVD data match, only relative trends can be observed. The calculation is slightly below the data at near-vertical directions and above it for intermediate zenith angles, except for a small patch in the SW direction ($\sim 225^{\circ}$ azimuth), where the situation is inverted. We calculate the $\chi^{2}/\mathrm{ndf}$ to be 1.34, where $\mathrm{ndf}$ is assumed to be the number of nonzero bins ($\mathrm{ndf}=31756$).
@@
In general, comparing the experimental and calculated angular distributions can reveal critical information about the accuracy of data analyses of underground experiments. The overall agreement obtained with the LVD experimental data suggests that mute can be a powerful and helpful tool to inform on potential sources of errors and misreconstructions in data analyses for underground experiments, mainly due to the absence of statistical errors.